# Honeybee: Decentralized Peer Sampling with Verifiable Random Walks for Blockchain Data Sharding

### v1

@@ 5.1.3. Adversary
For each of the three p2p networks, we simulate scenarios in which attackers control 5%, 10%, 20%, 30%, 40%, and 50% of the total nodes. The attacker-controlled dishonest nodes (Sybil nodes) are randomly chosen from all the nodes at the beginning of the simulation. To disturb the network and compromise potential victim node(s), the Sybil nodes employ strategies addressed in §5.1.4 during the simulation. Attackers have full control of their Sybil nodes and are able to build a desired topology with their Sybil nodes. We consider three types of initial Sybil node layouts as follows.
  • Mixed layout: the Sybil nodes are mixed into the p2p network. All the nodes in the network are connected to each other in a random way, and the Sybil nodes are simply among them. We consider this layout the most natural initial attacker layout.
  • Big cluster layout: the Sybil nodes form a big cluster. Most (98%) of the Sybil nodes only connect to each other while 2% of the Sybil nodes have connections to the honest cluster. The Sybil nodes whose connections include a connection to the honest nodes are named gateway Sybil nodes. The Sybil nodes who do not have a connection with the honest nodes are named trap Sybil nodes. The two types of Sybil nodes may switch their roles during the simulation.
  • Small clusters layout: the Sybil nodes form multiple small clusters. The Sybil nodes form 100 clusters each consists of roughly 81 nodes. Each Sybil cluster maintains one connection with the honest cluster. In other words, each Sybil cluster has one gateway Sybil nodes and the remaining nodes in the cluster are trap Sybil nodes.
We consider two types of victims: single victim node and multiple victim nodes. In the first case, Sybil nodes target a random honest node throughout the simulation (e.g., attackers attempt to compromise a client). In the second case, Sybil nodes target all honest nodes throughout the simulation (e.g., attackers attempt to compromise an organization or a company). In both cases, the goal of attackers is to eclipse the victim(s) by inserting as many Sybil nodes into the routing tables of the victim(s) as possible.
@@ 5.1.4. Adversary Strategies
Attackers can employ a wide range of adversary strategies using the nodes they control. Listed below are the strategies that we consider to be the most important. We categorize the strategies into two types: active strategies and passive strategies. We define the following active adversary strategies.
@@
  • Adversary routing: once a request reaches a dishonest node, the dishonest node may route the request in its favor to achieve certain goals. The goals include but are not limited to: preventing the request from reaching its intended destination, guiding the request to a dishonest destination, causing overhead (e.g., delays) for the request initiator, and causing inaccurate judgment (e.g., inaccurate neighbor scoring) for the request initiator.
  • Adversary peer selection: an honest node should select its peers in a random way or based on certain bona fide rules. An honest node should not make its peer selection decisions based on the identity of the candidate peers (and it should not be able to). However, a dishonest node may make such decisions based on the identity of the candidate peers. For example, a group of dishonest nodes may choose to add each other to their routing tables to form a Sybil cluster.
  • Equivocal table: a dishonest node can keep multiple copies of routing tables, each storing different peers, and use different routing tables to route different requests. This strategy can prevent the request from reaching its intended destination, guide the request to a dishonest destination, help other dishonest nodes with load-balancing, cause overhead (e.g., delays) for the request initiator, and cause inaccurate judgment (e.g., inaccurate neighbor scoring) for the request initiator.
We define the following passive adversary strategies.
@@
  • Adversary recommendation: upon request, an honest node should recommend (i.e., share the information of) a peer to another node in a random way or based on certain bona fide rules. An honest node should not make its recommendation decisions based on the identity of the request initiator and (and it should not be able to). However, a dishonest node may make such decisions based on the identity of the request initiator. For example, an honest victim’s dishonest neighbor may choose to recommend other dishonest nodes to the victim. This strategy is similar to request flood but in a passive manner.
  • Black hole: once a request reaches a dishonest node, the dishonest node may drop the request completely (i.e., being unresponsive to the request). This strategy can cause overhead (e.g., delays) and inaccurate judgment (e.g., inaccurate neighbor scoring) for the request initiator. In comparison with adversary routing, this strategy is less harmful but more difficult to detect/deter since honest nodes can also be unresponsive for various legitimate reasons.
@@ 5.2. Results
Figure 2. Nodes adhere to protocol: Sampling distribution from a random observation node in 100 thousand epochs. Every node follows its protocol. Node IDs are sorted by frequency in ascending order. True uniform sampling distribution is shown as a dashed line.
@@ 5.2.1. Sampling Distributions
Fig. 2a, 2b, and 2c show the sampling distributions for an arbitrarily chosen observation node with Honeybee , GossipSub, and Kademlia in a network of 16,384 honest nodes for 100 thousand epochs. In each setting, all nodes start (i.e., epoch 0) with random routing table configurations, and all nodes behave according to their protocol. In the figures, the true uniform sampling distribution (i.e., all nodes except for the observation node are sampled with equal probability) is shown as a dashed line. We observe that Honeybee  and the two baseline algorithms have similar sampling distributions when all nodes adhere to their protocol.
@@
(b) Chi-Square test results
(c) Chi-Square with 30% idle nodes
@@
To compare the sampling distributions of Honeybee  and the baseline algorithms with the true uniform sampling distribution, Fig. 3a plots the total variation distances between the three algorithms and the true uniform sampling distribution from the above experiment. The total variation distance curves suggest that sampling distributions from Honeybee  and the baseline algorithms converge to the true uniform sampling distribution throughout the 100 thousand epochs in a similar pattern. At the beginning of the simulations, the total variation distances from Honeybee  and the baseline algorithms are higher than 95%. At 100,000 epochs, the total variance distances from Honeybee  and the baseline algorithms decrease to roughly 23%. We also extended the simulations to 3 million epochs, and the total variation distances from the three algorithms decrease to 3-5%.
To examine the sampling distributions in detail, we conduct Chi-Square tests on the samples from Honeybee  and the baseline algorithms. We combine neighboring cells by dividing the node IDs into 127 node ID intervals of equal length (since the number of node IDs minus the observation node is divisible by 127). We divide the 100 thousand epochs into 10 time intervals of equal length. For each time interval, we examine the samples from that particular time interval. In Fig. 3b, we plot the Chi-Square values across the 10 time intervals for Honeybee  and the baseline algorithms. In Fig. 3c, we plot the Chi-Square values across the 10 time intervals for Honeybee  and the baseline algorithms when 30% of the nodes are idle (i.e., do not actively conduct sampling). We observe that, for Honeybee  and the baseline algorithms, we cannot reject the null hypothesis of uniform sampling with sufficient evidence when all nodes conduct sampling. When there are 30% of idle nodes, the Chi-Square values slightly increase for Honeybee  and GossipSub.
5.2.2. Why VRW and TCC
We demonstrate the importance of verifiable random walks and table consistency checks by showing the reader what happens when we do not use verifiable random walks or table consistency checks in Honeybee. Without the random walk verification mechanism, a random walk is not necessary random since dishonest nodes can hijack it by reconfigure their address tables according to their preferences. Similarly, without the table consistency check mechanism, dishonest nodes are able to store multiple copies of address tables and use different copies to route random walks according to their preferences. In other words, without verifiable random walks or table consistency checks, dishonest nodes can carry out a de facto adversarial routing attack.
(a) Honeybee  without VRW
(b) Honeybee  without TCC
@@ 5.2.3. Attack - Single Victim
Fig. 5a, 5b, and 5c show the ratio of dishonest peers in a random honest victim node’s routing table in 1,000 epochs for Honeybee  and the baseline algorithms in a network of 16,384 nodes, and 5%, 10%, and 20% of the total population consists of dishonest nodes. Fig. 6a, 6b, and 6c show the ratio of dishonest peers in a random honest victim node’s routing table in 1,000 epochs for Honeybee  and the baseline algorithms in a network of 16,384 nodes, and 30%, 40%, and 50% of the total population consists of dishonest nodes. Each node starts with a random routing table configurations. We observe that, under different levels of dishonest node percentage, Honeybee  outperforms the baseline algorithms in terms of the mean ratio of dishonest peers in the victim’s routing table.
Percent of dishonest nodes in the network
5% 10% 20% 30% 40% 50%
Algorithm Honeybee 6.16% 11.49% 21.91% 32.25% 40.90% 51.14%
GossipSub 9.99% 25.12% 42.24 55.52% 69.64% 81.79%
Kademlia 56.26% 63.22% 85.12% 87.15% 97.36% 95.23%
Table 1. Mean ratio of dishonest peers in victim’s routing table: Five honest nodes are randomly sampled as victims in five separate simulations with Honeybee, GossipSub, and Kademlia. They are attacked by 5%, 10%, 20%, 30%, 40%, and 50% of dishonest nodes.
To demonstrate that the results apply across different nodes in different simulations, we randomly sample five honest nodes as victims in five separate simulations. We calculate the mean ratio of dishonest peers in the victim’s routing table among the five nodes over the course of 1,000 epochs and present the results in Table. 1. We observe that Honeybee  consistently achieves near-uniform sampling and outperforms the baseline algorithms by 4-63% in terms of the mean ratio of dishonest peers in the victim’s routing table.
@@ 5.2.4. Attack - Multiple Victims
Fig. 7 shows the cumulative number of honest nodes that can be eclipsed by dishonest nodes in a network with 50% dishonest nodes in 1,000 epochs with Honeybee, GossipSub, and Kademlia. There are 8,192 honest nodes in total. We observe that, in 1,000 epochs, dishonest nodes can eclipse over 75% of honest nodes with GossipSub and Kademlia while they cannot eclipse honest nodes with Honeybee.

### v4

@@ 4.1.3. Adversary
For each of the three p2p networks, we simulate scenarios in which attackers control 5%, 10%, 20%, 30%, 40%, 50%, 60%, 70%, and 80% of the total nodes. The attacker-controlled dishonest nodes (e.g., Sybil nodes) are randomly chosen from all the nodes at the beginning of the simulation. To disturb the network and compromise potential victim node(s), the dishonest nodes employ strategies addressed in §4.1.4 during the simulation. Attackers have full control of their dishonest nodes and are able to build a desired topology with their dishonest nodes. We consider two types of initial dishonest node layouts: mixed layout and clustered layout.
  • Mixed layout: the dishonest nodes are mixed into the p2p network. All the nodes in the network are connected to each other in a random way, and the dishonest nodes are simply among them. We consider this layout the most natural initial attacker layout.
  • Clustered layout: the dishonest nodes form one or more clusters. Most ($\approx$98%) of the dishonest nodes only connect to each other while 2% of the dishonest nodes have connections to the honest cluster. The dishonest nodes whose connections include a connection to the honest nodes are named gateway Sybil nodes. The dishonest nodes who do not have a connection with the honest nodes are named trap Sybil nodes. The two types of dishonest nodes may switch their roles during the simulation.
Figure 2. In a Honeybee network with 50% dishonest nodes, all dishonest nodes cluster to attack a victim whose initial address table contains only one honest outgoing neighbor (an unlikely adverse scenario). We conduct five separate such experiments and present the worst-case outcome for the victim.
At first glance, a lethal attack in the clustered layout setting in which an honest victim node whose random walks venture into the dishonest nodes’ cluster is completely trapped (eclipsed) within the dishonest node cluster appears feasible. However, in Fig. 2, we show that such an attack is unlikely to succeed if the number of dishonest nodes is not significantly greater than the number of honest nodes (the settings considered in our experiments). If the victim node has even a single outgoing neighbor in the honest cluster (i.e., is not totally eclipsed), incoming connections requests from the honest cluster quickly increase the fraction of honest peers within the victim’s address table. Moreover, with verifiable random walks, attracting a target victim node’s walks inside the adversary’s cluster is also not easy. In our evaluations, we therefore focus on the mixed layout as the more damaging setting to the honest nodes.
We focus on two types of victims: single victim node and multiple victim nodes. In the first case, dishonest nodes target a random honest node throughout the simulation (e.g., attackers attempt to compromise a client). In the second case, dishonest nodes target all honest nodes throughout the simulation (e.g., attackers attempt to compromise an organization or a company). In both cases, the goal of attackers is to eclipse the victim(s) by inserting as many dishonest nodes into the address tables of the victim(s) as possible. [^7] In this paper, an eclipse refers to a full eclipse—a situation where all the neighbors of one or more victims are dishonest.
@@ 4.1.4. Adversary Strategies
In our evaluation, attackers employ a wide range of adversary strategies using the nodes they control. We categorize the adversary strategies into two types: active strategies and passive strategies. The active adversary strategies include request flood, adversarial routing, adversarial peer selection, and equivocal table defined as follows:
@@
  • Adversarial routing: once a request reaches a dishonest node, the dishonest node may route the request in its favor to achieve certain goals. The goals include but are not limited to: preventing the request from reaching its intended destination, guiding the request to a dishonest destination, causing overhead (e.g., delays) for the request initiator, and causing inaccurate judgment (e.g., inaccurate neighbor scoring) for the request initiator.
  • Adversarial peer selection: an honest node should select its peers in a random way or based on certain bona fide rules. An honest node should not make its peer selection decisions based on the identity of the candidate peers (and it should not be able to). However, a dishonest node may make such decisions based on the identity of the candidate peers. For example, a group of dishonest nodes may choose to add each other to their routing tables to form a dishonest cluster.
  • Equivocal table: a dishonest node can keep multiple copies of routing tables, each storing different peers, and use different routing tables to route different requests. This strategy can prevent the request from reaching its intended destination, guide the request to a dishonest destination, help other dishonest nodes with load-balancing, cause overhead (e.g., delays) for the request initiator, and cause inaccurate judgment (e.g., inaccurate neighbor scoring) for the request initiator. Equivocal table attacks are dangerous because they are important means to further attacks.
The passive adversary strategies include selective request acceptance, adversarial recommendation, and black hole defined as follows:
@@
  • Adversarial recommendation: upon request, an honest node should recommend (i.e., share the information of) a peer to another node in a random way or based on certain bona fide rules. An honest node should not make its recommendation decisions based on the identity of the request initiator and (and it should not be able to). However, a dishonest node may make such decisions based on the identity of the request initiator. For example, an honest victim’s dishonest neighbor may choose to recommend other dishonest nodes to the victim. This strategy is similar to request flood but in a passive manner.
  • Black hole: once a victim request reaches a dishonest node, the dishonest node may ignore the request completely (i.e., being unresponsive to the request). This strategy can cause overhead (e.g., delays) and inaccurate judgment (e.g., inaccurate neighbor scoring) for the request initiator. In comparison with adversarial routing, this strategy is less harmful because an honest node can simply drop its random walk. Modern p2p networks use reputation or credit systems to help mitigate black hole attacks  (Stepień and Poniszewska-Marańda, 2020; Dixit et al., 2015).
@@ 4.2. Results
Figure 3. Nodes adhere to protocol: Sampling distribution from a random observation node in 100,000 epochs. Every node follows its protocol. Node IDs are sorted by frequency in ascending order. True uniform sampling distribution is shown as a dashed line.
@@ 4.2.1. Sampling Distributions
Fig. 3a, 3b, and 3c show the sampling distributions for an arbitrarily chosen observation node with Honeybee, GossipSub, and Kademlia in a network of 16,384 honest nodes for 100,000 epochs. In each setting, all nodes start (i.e., epoch 0) with random routing table configurations, and all nodes behave according to their protocol. In the figures, the true uniform sampling distribution (i.e., all nodes except for the observation node are sampled with equal probability) is shown as a dashed line. We observe that Honeybee and the two baseline algorithms have similar sampling distributions when all nodes adhere to their protocol.
@@
(b) $\chi^{2}$ test results
(c) $\chi^{2}$ with 30% idle nodes
@@
To compare the sampling distributions of Honeybee and the baseline algorithms with the true uniform sampling distribution, Fig. 4a plots the total variation distances (TVDs) between the three algorithms and the true uniform sampling distribution from the above experiment. The TVDs suggest that sampling distributions from Honeybee and the baseline algorithms converge to the true uniform sampling distribution throughout the 100,000 epochs in a similar pattern. At the beginning of the simulations, the TVDs from Honeybee and the baseline algorithms are higher than 95%. At 100,000 epochs, the total variance distances from Honeybee and the baseline algorithms decrease to roughly 23%. We also extended the simulations to 3 million epochs, and the TVDs from the three algorithms decrease to 3-5%.
To examine the sampling distributions in detail, we conduct Chi-Square tests on the samples from Honeybee and the baseline algorithms. We combine neighboring cells by dividing the node IDs into 127 node ID intervals of equal length (since the number of node IDs minus the observation node is divisible by 127). We divide the 100,000 epochs into 10 time intervals of equal length. For each time interval, we examine the samples from that particular time interval. In Fig. 4b, we plot the Chi-Square values across the 10 time intervals for Honeybee and the baseline algorithms. In Fig. 4c, we plot the Chi-Square values across the 10 time intervals for Honeybee and the baseline algorithms when 30% of the nodes are idle (i.e., do not actively conduct sampling). We observe that, for Honeybee and the baseline algorithms, we cannot reject the null hypothesis of uniform sampling with sufficient evidence when all nodes conduct sampling. When there are 30% of idle nodes, the Chi-Square values increase slightly for Honeybee and GossipSub. The takeaway here is that Honeybee and the baseline algorithms can all conduct near-uniform peer sampling in the absence of adversarial nodes.
@@
(a) 60% dishonest nodes
(b) 70% dishonest nodes
(c) 80% dishonest nodes
Figure 7. An honest node attacked by 60%, 70%, and 80% of dishonest nodes: Single random honest node under attack in Honeybee, GossipSub, and Kademlia.
@@ 4.2.2. Attack - Single Victim
Fig. 5, Fig. 6, and Fig. 7 show the ratio of dishonest peers in a random honest victim node’s routing table in 1,000 epochs for Honeybee and the baseline algorithms in a network of 16,384 nodes, and 5%, 10%, 20%, 30%, 40%, 50%, 60%, 70%, and 80% of the total population consists of dishonest nodes. The goal of all the dishonest nodes is to eclipse the victim node. All nodes start with random routing table configurations. We observe that, under different levels of dishonest node percentage, Honeybee consistently outperforms the baseline algorithms in terms of the mean ratio of dishonest peers in the victim’s routing table.
Percentage of dishonest nodes in the network
5% 10% 20% 30% 40% 50% 60% 70% 80%
Honeybee 6.16% 11.49% 21.91% 32.25% 40.90% 51.14% 61.44% 71.73% 94.62%
GossipSub 9.99% 25.12% 42.24 55.52% 69.64% 81.79% 91.64% 96.26% 99.49%
Kademlia 56.26% 63.22% 85.12% 87.15% 97.36% 95.23% 97.73% 98.80% 98.64%
Table 1. Mean ratio of dishonest peers in victim’s routing table over 1,000 epochs: Five honest nodes are randomly sampled as victims in five separate simulations with Honeybee, GossipSub, and Kademlia. We consider different percentage of dishonest nodes in the entire network ranging from 5% to 80%.
To demonstrate that the results apply across different nodes in different simulations, we randomly sample five honest nodes as victims in five separate simulations. We calculate the mean ratio of dishonest peers in the victim’s routing table among the five nodes over the course of 1,000 epochs and present the results in Table. 1. We observe that Honeybee consistently achieves near-uniform sampling and outperforms the baseline algorithms by 4-63% in terms of the mean ratio of dishonest peers in the victim’s routing table when 5-70% of the nodes in the network are dishonest. While Honeybee’s performance drops when 80% of the nodes in the network are dishonest (a scenario unlikely in a usable real-world p2p network), it still outperforms the baseline algorithms by 4%.
@@ 4.2.3. Attack - Multiple Victims
Fig. 8 shows the cumulative number of honest nodes that can be eclipsed by dishonest nodes in a network with 50% dishonest nodes in 1,000 epochs with Honeybee, GossipSub, and Kademlia. There are 8,192 honest nodes in total, and their goal is to eclipse (or partition, since we are addressing multiple victim nodes) as many honest nodes as possible. We observe that, in 1,000 epochs, dishonest nodes can eclipse over 75% of honest nodes with GossipSub and Kademlia while they cannot eclipse honest nodes with Honeybee.
(a) Initialized with 62.5% of Sybils
(b) Initialized with 75% of Sybils
(c) Initialized with 87.5% of Sybils
Figure 9. An honest Honeybee node attacked by 50% of dishonest nodes in the network: when the honest node’s initial routing table happens to be poorly configured.
4.2.4. Bad Start
We have discussed how a Honeybee node bootstraps its initial routing table configuration in §2. It is possible for an honest Honeybee node to start off with a poorly configured routing table (i.e., the ratio of dishonest nodes in one’s initial routing table configuration is significantly higher than that of the entire network). To see what happens when nodes get a bad initial routing table configuration, we run experiments on Honeybee networks with 50% of dishonest nodes. In these experiments, we identified random, unfortunate Honeybee nodes whose initial routing tables contain 62.5%, 75%, and 87.5% of dishonest nodes. In each experiment, all the dishonest nodes in the network target and attempt to eclipse the honest node under our observation. We then examined their peer sampling performances under poorly initialized routing tables. The results are shown in Fig. 9. We observe that, even with a poorly initialized routing table, Honeybee nodes can still conduct near-uniform sampling, and they converge to $\epsilon$-uniform sampling with $\epsilon=0.03$ quickly (within 50 epochs). Therefore, we conclude that poorly initialized routing tables do not adversely affect Honeybee nodes’ peer sampling performance.
4.2.5. Why VRW and TCC
(a) Honeybee without VRW
(b) Honeybee without TCC
We demonstrate the importance of verifiable random walks and table consistency checks by showing the reader what would happen if we did not use verifiable random walks or table consistency checks in Honeybee. Without the random walk verification mechanism, a random walk is not necessary random since dishonest nodes can hijack it by reconfigure their address tables according to their preferences. Similarly, without the table consistency check mechanism, dishonest nodes are able to store multiple copies of address tables and use different copies to route random walks according to their preferences. In other words, without verifiable random walks or table consistency checks, dishonest nodes can carry out a de facto adversarial routing attack.
4.3. Practical Considerations
So far, we have demonstrated Honeybee’s effectiveness and efficiency with analyses and experiments. Now, we would like to briefly discuss some practical considerations when it comes to applying Honeybee in real-world networks.
In §2, we have mentioned that Honeybee can use block headers as its source of randomness. However, what if the block headers are not secure (e.g., they may be susceptible to tampering, such as grinding attacks (Hackn.gitbook.io, 2025)) in the blockchain network we are working on? To address this issue, we can introduce an extra layer of protection to the consensus mechanism using a verifiable delay function (VDF) (Boneh et al., 2018), which ensure that randomness is both secure and verifiable. Moreover, we may want to apply Honeybee to general p2p networks which do not have a blockchain. As we have discussed in §3, Honeybee needs a shared source of randomness for nodes to perform decentralized peer sampling. This source of randomness does not necessarily have to be derived from a blockchain, as long as it is secure and resistant to manipulation. A good example of such a secure source of randomness is decentralized randomness beacons (Cloudflare, 2025), which are designed to generate unbiased and publicly verifiable randomness in a decentralized manner. By incorporating a reliable shared source of randomness, we can extend the applicability of Honeybee beyond blockchain-based systems to a broader range of p2p networks.
Note that the frequency of shared randomness generation does not impact the frequency or time consumption of random walks and peer samplings. The reasons are as follows: (1) Popular blockchains (e.g., Ethereum (Etherscan, 2024a)) and decentralized randomness beacons (e.g., Drand (Drand.love, 2024)) are able to provide shared randomness within tens of seconds. (2) As mentioned in §3.3, in each round of Honeybee, a fraction $\eta$ of nodes become eligible and perform peer sampling. Even if the time required to generate fresh random values is a lot longer (e.g., at least twice as long, which is uncommon) than the typical time required to perform random walks, we can generate extra pseudo-random values using amplification (Colbeck and Renner, 2012). By incorporating an amplification process, more than one round of peer sampling can be conducted.
As for communication overhead, in a Honeybee network with $n$ nodes, a random walk consists of at most $O(log(n))$ hops, which correspond to $O(log(n))$ messages. Suppose a node’s routing table size is $k$, a node only needs to exchange $O(klog(n))$ messages to update its entire routing table. The communication overhead for all nodes in the network to update their entire routing tables is just $O(knlog(n))$. Regarding the time required for all nodes to update their entire routing tables, since in each round we allow a fraction $\eta$ of nodes to perform peer sampling, the network requires $\frac{k}{2\eta}$ rounds for all nodes to update their entire routing tables. To calculate the wall-clock time, assuming $k=24$, $\eta=0.1$ and a round costs 12 seconds (Etherscan, 2024a), the total time consumption would be $24$ minutes. Even if we relax the assumption by setting $\eta=0.001$ (i.e, only $0.1\%$ of nodes can conduct peer sampling in a round), the network would still allow all nodes to update their entire routing tables within $1.67$ days, which should be more than acceptable in most p2p network applications. Further improving the efficiency of Honeybee peer sampling is an interesting direction for future work.
In terms of adoption, networks employing other p2p protocols can smoothly transition into Honeybee in a hybrid manner. During the transition, the network encourages nodes to adapt to Honeybee while allowing the use of other protocols. Honeybee nodes work normally with each other as we described in §3. In addition, Honeybee nodes store the addresses of non-Honeybee nodes in a separate table to stay in touch with them. Non-Honeybee nodes can choose to adapt to Honeybee at their own discretion.